### v1

# Multi-Modal Language Models as Text-to-Image Model Evaluators

Abstract
TThe steady improvements of text-to-image (T2I) generative models lead to slow deprecation of automatic evaluation benchmarks that rely on static datasets, motivating researchers to seek alternative ways to evaluate the T2I progress. In this paper, we explore the potential of multi-modal large language models (MLLMs) as evaluator agents that interact with a T2I model, with the objective of assessing prompt-generation consistency and image aesthetics. We present Multimodal Text-to-Image Eval (MT2IE), an evaluation framework that iteratively generates prompts for evaluation, scores generated images and matches T2I evaluation of existing benchmarks with a fraction of the prompts used in existing static benchmarks. Moreover, we show that MT2IE’s prompt-generation consistency scores have higher correlation with human judgment than scores previously introduced in the literature. MT2IEigenerates prompts that are efficient at probing T2I model performance, producing the same relative T2I model rankings as existing benchmarks while evaluating on 80×80$\times$80 × less prompts.
[…]
1 Introduction
With the ever-growing use of generated images in academic and commercial settings (Rombach et al., 2022; Podell et al., 2023; Sinha et al., 2024), the evaluation of text-to-image (T2I) generative models is a rapidly growing and highly relevant area of research. Aesthetic quality, diversity of generated samples, and image-text consistency are crucial for evaluating T2I models and driving the development of reliable, high-quality image generators.
Image-text consistency in particular has gained more attention as users’ text prompts become longer and more semantically complex. For prompt consistency, there is no consensus on a single best evaluation benchmark. Generally, static sets of text prompts are used to benchmark model progress. These prompts can be human-written, typically sourced from image captioning datasets (e.g., COCO (Lin et al., 2014), Conceptual Captions (Changpinyo et al., 2021)). Other benchmarks are automatically generated from free-form text, and designed specifically for image generation (e.g., ImagenHub (Ku et al., 2023), ConceptMix (Wu et al., 2024)).
Figure 1: We present Multimodal Text-to-Image Eval (MT2IE), our novel evaluation methodology for text-to-image (T2I) generation. Previous approaches to T2I model evaluation require large, static benchmarks and often multiple external models. Our approach in a single system that both generates and scores prompts. We show that MT2IE  can generate as few as 20 prompts while still maintaining consistent model rankings of other approaches that use more compute and more data.
Evaluation frameworks that use static prompt datasets most commonly use automatic metrics that assess how well a generated image matches a given text prompt. There are alignment-based metrics, which typically judge the quality of the generated image by computing similarity between the image and prompt. CLIPScore (Hessel et al., 2021) is a widely used alignment metric that leverages the CLIP model (Radford et al., 2021) to calculate cosine similarity between generated images and text prompts. However, its effectiveness is limited by the weaknesses of CLIP, which struggles with capturing compositional vision-language aspects such as spatial relationships, object cardinality, and attribute binding (Tong et al., 2024; Thrush et al., 2022b; Yuksekgonul et al., 2023). Many generated question-answering-based (GQA) metrics  (Hu et al., 2023; Cho et al., 2024; Lin et al., 2024) have been introduced to address the issues of fine-grained T2I evaluation. These use LLMs to generate question sets about the text prompt and then separate VQA models to score the generated questions given the prompt’s generated image.
Recently, multimodal LLMs (MLLMs) (Alayrac et al., 2022; Liu et al., 2023) have been developed to enhance large language models (LLMs) with vision understanding capabilities, resulting in state-of-the-art performance across some vision-language tasks. Various evaluation approaches use MLLMs for prompt consistency evaluation, often via some form of visual question answering (Lu et al., 2023; Ku et al., 2024; Zhang et al., 2023; Meng et al., 2024; Tan et al., 2024). Although these approaches leverage MLLMs for improved evaluation, they still require fixed prompt datasets without any ability for dynamic evaluation. This can create a narrow focus on optimizing T2I performance for predefined prompts, rather than driving broad improvements. Additionally, rapid advancements in model capabilities can render static prompt sets ineffective, as they become overly simplistic and fail to provide meaningful evaluation (Corry et al., 2021; Ott et al., 2022; Gupta et al., 2024). Across benchmarks and metrics, the use of static evaluation frameworks is prone to overfitting and vulnerable to deprecation.
In this work, we present Multimodal Text-to-Image Eval (MT2IE), a unified framework that moves beyond static evaluation by iteratively generating dynamic prompts and scoring image generations. We move beyond the use of fixed prompt sets by exploring the potential of MLLMs as evaluator agents that interact with a T2I model. Instead of relying on a static benchmark of prompts for image generation, the MLLM is now responsible for both generating text prompts and scoring the consistency of the generated images based on the prompts. This removes the need to curate a prompt set for benchmarking while performing end-to-end evaluation with a single model. This additionally allows us to leverage the MLLM to generate increasingly complex prompts for generation, pushing the limits of the T2I model, and adaptive prompts, automatically rewriting prompts according to the T2I model’s performance.
In our approach, we first focus on validating the capabilities of off-the-shelf MLLMs to generate questions and score prompt-generation consistency of T2I models, showing that a single model can outperform existing human judgment correlation metrics on the static TIFA160 benchmark (Hu et al., 2023). This contrasts with previous work that either relied on proprietary GPT-4 variants (Zhang et al., 2023) or found open-source MLLMs’ judgments ineffective and uncorrelated with human judgments (Ku et al., 2024; Jiang et al., 2024). Second, we leverage MLLMs to not only score but also generate prompts of increasing complexity to evaluate T2I models. We use this MLLM-based approach to compare the prompt-generation consistency of 8 state-of-the-art T2I models and find that the rankings from our approach are comparable to those reported by static benchmarks, while only using a fraction ($\sim$1%) of prompts. We also show that MT2IE is easily generalizable to evaluate aspects of image generation beyond prompt-generation consistency, with results for aesthetic quality presented in Appendix appendix A. Finally, we investigate MLLMs as evaluator agents for prompt-generation consistency by enabling the interaction between the MLLM and the T2I model.
Our contributions are as follows:
  1. 1.
    We introduce Multimodal Text-to-Image Eval (MT2IE), an evaluation framework that uses MLLMs as evaluator agents that interact with T2I models. MT2IE enables open-source MLLMs to be effective T2I model evaluators, contrasting with previous work that relied on proprietary GPT-4 variants (Zhang et al., 2023) or found open-source MLLMs’ judgments ineffective and uncorrelated with human judgments (Ku et al., 2024; Jiang et al., 2024). To our knowledge, we are the first to show that open-source MLLMs can be effective end-to-end T2I model evaluators.
  2. 2.
    We show MLLMs can rank the prompt-generation consistency and image aesthetics of T2I models comparably to static benchmarks, while only requiring a fraction of prompts. These prompts are generated by the same MLLM.
  3. 3.
    We observe MT2IE exhibits better correlation with human judgements than existing metrics on the static benchmarks such as TIFA and DSG.
  4. 4.
    We demonstrate that MT2IE can also adaptively adjust prompts based on T2I model performance on previous prompts, allowing for bespoke model evaluation.
2 Related Work
Alignment-based metrics.
Alignment-based metrics use similarity scores between image and text in the embedding space of a vision-language model (VLM). VLMs such as CLIP (Radford et al., 2021) or BLIP (Li et al., 2022) are commonly used as an automatic image-text alignment metric. The most long-standing of these metrics is CLIPScore (Hessel et al., 2021), which computes the cosine similarity of the image and text CLIP embeddings. However, many works have shown that CLIP and other VLMs treat text as a bag-of-words, failing to capture the semantics of compositional text that includes multiple objects and their attributes and spatial relations (Yuksekgonul et al., 2023; Ma et al., 2023; Thrush et al., 2022b; Kamath et al., 2023). Thus their similarity estimates are not reliable, especially when text prompts specify cardinality or compositional reasoning (Wang et al., 2023; Ma et al., 2023; Kamath et al., 2023). Despite this, many recent, impactful image and video generation methods still use CLIPScore as a prompt consistency metric (Ruiz et al., 2023; Brooks et al., 2023; Chang et al., 2023; Meng et al., 2022; Kang et al., 2023; Ho et al., 2022; Wu et al., 2023a). Similarity-based approaches have been enhanced by collecting large scale human ratings of generated images to enable finetuning of CLIP (Kirstain et al., 2023; Wu et al., 2023b). Similarly, human preference data has been used to create reward models out of alternative VLMs such as BLIP (Xu et al., 2023). Of these methods, it is unclear if any have achieved as widespread adoption as CLIPScore.
GQA-based metrics.
Generated question answering (GQA) metrics for image-text consistency were introduced to overcome some of the above-described limitations. These metrics use a language model to generate questions and a VQA model to score the image using the generated questions (Cho et al., 2024; Hu et al., 2023; Changpinyo et al., 2022; Yarom et al., 2023). Interestingly, while more fine-grained, these metrics are not necessarily more reliable than alignment-based metrics (Saxon et al., 2024; Ross et al., 2024). For instance, a strong yes-bias is exhibited by VQA and language models (Goyal et al., 2017; Ross et al., 2024), which are used to generate and score questions.
Large language models as evaluators.
Following the use of LLMs for image-prompt alignment question generation, LLMs or multimodal LLMs are also being used as standalone evaluators (Lu et al., 2023; Ku et al., 2024; Zhang et al., 2023; Tan et al., 2024). Lu et al.  (Lu et al., 2023) use a VLM to generate a text description of a generated image, then asks an LLM to judge the faithfulness of the image description and the image prompt. Ku et al. (Ku et al., 2024) and Zhang et al. (Zhang et al., 2023) prompt MLLMs for image-prompt faithfulness and overall image quality scores. Both works focus on using proprietary, multimodal versions of GPT-4, finding that the MLLM used has great effect on correlation with human evaluators. Notably, Ku et al. (Ku et al., 2024) find that open-sourced MLLMs are significantly worse at generated image evaluation, with Llava (Liu et al., 2023) having 90% worse human correlation compared to GPT-4v when using their evaluation methodology.
3 Experimental Overview
We demonstrate the effectiveness of using MLLMs as evaluator agents for assessing T2I prompt-generation consistency. In Multimodal Text-to-Image Eval (MT2IE), MLLM interacts with the T2I model by proposing text prompts and scoring resulting generations. We show that our interactive approach enables MLLMs to effectively benchmark T2I models in two distinct settings: either by systematically generating prompts of increasing complexity or by dynamically adapting prompts based on T2I model performance. In the following sections, we focus on evaluating image-text consistency as it directly impacts the overall visual quality of generated images (Astolfi et al., 2024; Xu et al., 2023; Kirstain et al., 2023). Additionally, in Appendix appendix A we show that MT2IE is an easily generalizable framework, capable of evaluating other aspects of T2I models such as image aesthetics. We use MT2IE to evaluate a total of 8 recent T2I models including latent diffusion model (LDM) v2.1 (Rombach et al., 2022), LDM v3 (Esser et al., 2024), LDM XL (Podell et al., 2023), LDM XL-Turbo (Sauer et al., 2025), Flux (Labs, 2024), a cascaded pixel-based diffusion model (CDM) (Raffel et al., 2020), Playground 2.5 (Li et al., 2024b) and Pixart-$\alpha$ (Chen et al., 2023). We provide an outline of our experiments and elaborate in subsequent sections.
Experiment 1: Validation of MLLM prompt consistency scoring capabilities.
We validate that open-sourced MLLMs are capable of scoring image-text consistency on existing benchmarks without the usage of other models or fine-tuning. Existing works that use MLLMs for end-to-end evaluation either do not use existing benchmarks or fine-tune open-sourced MLLMs for the task (Tan et al., 2024; Ku et al., 2024). Existing GQA-based benchmarks rely on a system of models for question generation and scoring; typically an LLM is used to generate questions based on a curated text prompt, the generated questions are validated by an additional LLM or question-answering model, then finally a VQA model scores the image using the generated questions (Cho et al., 2024; Yarom et al., 2023; Hu et al., 2023; Gupta and Kembhavi, 2022). We validate that MLLMs are capable of emulating GQA-based benchmarks by generating questions about a given prompt and scoring the generated questions given the corresponding image. We evaluate using TIFA160 (Hu et al., 2023), and show that a single open-sourced MLLM can recreate this benchmark and score questions with better correlation to human judgment. We detail this experiment in Section section 4.
Experiment 2: MLLMs to generate prompt consistency benchmarks.
After validating MLLMs’ abilities to generate and score prompt consistency questions in Experiment 1, we then introduce MT2IE , demonstrating how MLLMs can iteratively generate a full T2I benchmark and judge model performance. In particular, we show that MLLM-generated benchmarks preserve relative model rankings computed by human annotators in prior work. We also show that our method is efficient, accurately ranking models with significantly fewer prompts than existing benchmarks of handwritten prompts. We detail these experiments in Section section 5.
Experiment 3: MLLMs to generate interactive prompt consistency benchmarks.
Beyond just generating benchmarks, we demonstrate that MT2IE can act as an evaluator agent, adapting to model-specific performance. Specifically, we show that MLLMs can adjust the complexity of prompts based on model performance on previous prompts. This enables each individual T2I model to have bespoke prompts tailored to its abilities, discussed in Section section 6.
4 Experiment 1: Validation on Static Benchmarks & Metrics
We first validate that off-the-shelf, open-source MLLMs can effectively reproduce the results of widely used scoring approaches on benchmarks of static prompts. We explore 3 different MLLMs’ abilities to generate and score questions: Molmo-7B (Deitke et al., 2024), Llama-3.2VI-11B (AI, 2024), and Llava-v1.6-34B (Liu et al., 2023). We selected only open-source models, as existing work on MLLMs as T2I evaluators focus on using proprietary models (Ku et al., 2024; Zhang et al., 2023), and we want to demonstrate that open-source MLLMs work well in our framework. We evaluate MLLMs’ ability to recreate the TIFA160 benchmark (Hu et al., 2023) and use a singe MLLM to execute two different scoring approaches: (i) a generated answering (GQA) approach, where questions about a prompt are generated by an LLM and answered by a separate VQA model (e.g., (Hu et al., 2023; Cho et al., 2024; Yarom et al., 2023; Changpinyo et al., 2022; Cho et al., 2023)) and (ii) a VQAScore based approach (Lin et al., 2024), which computes the likelihood of answering “Yes” to the question “Does this figure show {text prompt}. Please answer yes or no.".
4.1 Generated Question Answering Approach
GQA-based approaches use an LLM to generate questions based on prompts then score generated questions and images using a VQA model. The performance of the VQA model being used is crucial as VQA accuracy has direct correlation with human judgment (Hu et al., 2023). For this reason, we first confirm that MLLMs have strong VQA performance on existing GQA benchmarks by using the TIFA160 LLM-generated question set. We compare the MLLMs’ VQA accuracy to the 2 best VQA models from Hu et al. (2023); results are in Figure figure 9 (Appendix appendix C) and show that all 3 MLLMs outperform the VQA models used in existing GQA benchmarks.
Table 1: Human judgment correlations of existing VQA-based T2I consistency benchmarks and MT2IE, measured by Spearman’s $\rho$ and Kendall’s $\tau$ on TIFA160. End-to-end MLLMs have higher correlation with human judgements than existing GQA benchmarks using LLMs and VQA models.
$\rho$ $\tau$
Existing GQA Benchmarks
TIFA (Instruct-BLIP) 46.0 36.0
VQ2A (Instruct-BLIP) 42.6 35.2
DSG (PaLI) 57.1 45.8
MLLM Question Generation and Scoring
Llama-3.2-11B-Vision-Instruct 60.8 46.3
Molmo-7b-d 59.0 45.9
Llava-v1.6-34b 61.3 47.6
With this validation, we now investigate if a single MLLM can effectively generate and score questions given a text prompt and generated image. This constitutes a benchmark generation task, where the MLLM is responsible for creating new questions and scoring them. By consolidating these tasks to one model, this approach eliminates the resource-intensive reliance on external, often proprietary LLMs. We use all 3 MLLMs to generate and score questions about all prompts in TIFA160. The system prompts used to generate questions is shown in Appendix appendix C.
To assess quality of the generated questions, we follow previous work and measure the correlation between the VQA accuracy (from LLM-generated questions from existing benchmarks (Cho et al., 2024; Hu et al., 2023; Gupta and Kembhavi, 2022) or MLLM-generated questions in our work) and human ratings of image-text consistency using a 1-5 Likert scale. We use images and human ratings from prior work (Cho et al., 2024) on 5 different T2I models: LDM1.1 (Rombach et al., 2022), LDM1.5 (Rombach et al., 2022), LDM2.1 (Rombach et al., 2022), minDALL-E (Saehoon et al., 2021), and VQ-Diffusion (Gu et al., 2022). We show the Spearman’s $\rho$ and Kendall’s $\tau$ in Table table 1.
Notably, every MLLM achieves higher correlation with human annotations than existing GQA-based approaches. These results show that open-sourced MLLMs can effectively execute all tasks performed by separate models in existing GQA benchmarks. We additionally compare the GPT3-generated questions in TIFA160 and the MLLM-generated questions from Llava-v1.6-34b, the LLM with the highest human correlations. We analyze question statistics and BLEU score (Papineni et al., 2002) (a commonly used NLP metric to compare two sets of text) and find that entities in both sets of questions are very similar. Detailed results on generated question set verification are in Appendix appendix G.
4.2 VQAScore Approach
Table 2: Human judgment correlations of existing automatic T2I consistency metrics, including VQAScore Lin et al. (2024) computed by the CLIP-FlanT5 model and VQAScore computed by zero-shot MLLMs; measured by Spearman’s $\rho$ and Kendall’s $\tau$ on TIFA160.
$\rho$ $\tau$
Automatic Metrics
BLEU-4 18.3 18.2
ROUGE-L 32.4 24.0
METEOR 34.0 27.4
SPICE 32.8 23.2
CLIPScore (ViT-B-32) 28.7 19.1
MLLM VQAScore
VQAScore (CLIP-FlanT5, previous SOTA) 66.9 52.8
VQAScore (Llama-3.2-11B-Vision-Instruct) 52.1 43.7
VQAScore (Molmo-7b-d) 49.6 38.4
VQAScore (Llava-v1.6-34b) 67.5 54.4
VQAScore (Lin et al., 2024) computes the likelihood of the model generating “yes" when asked “Does this figure show {text prompt}. Please answer yes or no.". VQAScore has been shown to be more reliable than GQA-based approaches and alignment metrics like CLIPScore, likely due to CLIPScore’s challenges with object composition  (Lin et al., 2024; Tong et al., 2024; Thrush et al., 2022a; Yuksekgonul et al., 2023), and the tendency of GQA-based approaches to exhibit a “yes-bias" in question answering tasks (Ross et al., 2024). We compare the correlations with human ratings on TIFA160 between VQAScore calculated using the 3 aforementioned MLLMs, VQAScore calculated with the CLIP-FlanT5 model (which was fine-tuned for SOTA VQAScore calculation (Lin et al., 2024)), and CLIPScore. We use the same human ratings as in Section section 4.1; results are in Table table 2. All MLLMs outperform CLIPScore in human correlation, but only Llava-v1.6-34b achieves higher correlation with human ratings than CLIP-FlanT5.
We find that open-sourced MLLMs can replace systems of LLM and VQA models commonly used in existing QGA benchmarks. Llava-v1.6-34b achieved the best correlation with human ratings when used to compute VQAScore as prompt consistency scores; so we use Llava-v1.6-34b with VQAScore as our MLLM evaluator for further experiments. Experiments with question generation and answering for prompt consistency scoring are in Appendix appendix C. This validation lays the framework for our experiments, demonstrating that open-sourced MLLMs can be used for end-to-end T2I evaluation.
5 Experiment 2: MLLMs for Benchmark Generation
iteration 1: a dog running through a park, chasing a ball Consistency Score: 0.977 Refer to caption iteration 2: a dog running through a park chasing a ball with a frisbee in the air
Consistency Score: 0.938
Refer to caption iteration 3: a dog running through a park chasing a ball with a frisbee in the air with a bird flying overhead
Consistency Score: 0.861
Refer to caption iteration 4: a dog running through a park chasing a ball with a frisbee in the air with a bird flying overhead and a person walking on a path in the background
Consistency Score: 0.944
Refer to caption iteration 5: a dog running through a park chasing a ball with a frisbee in the air a bird flying overhead and a person walking on a path in the background with a small wooden bench on the side of the path
Consistency Score: 0.929
Figure 2: An example our method MT2IE generating progressively difficult prompts, with additions shown in purple. Corresponding generated images and scores are also shown. Prompts become more complex while maintaining linguistic structure.
(a) The progression of model scores as the prompts become increasingly more difficult.
(b) Shifts in various linguistic properties, averaged across all prompts in a given iteration.
Figure 3: Results for MT2IE, where the MLLM iteratively generates progressively more difficult prompts over 5 iterations.
In Section section 4, we showed that MLLMs are able to reproduce static benchmarks using both question-answering and prompt likelihood approaches. We now introduce our approach, the MT2IE framework, and show that it can be used to iteratively generate new, efficient benchmarks for image-text consistency. With each iteration, the generated prompts increase in complexity and thus challenge the T2I models in progressively more difficult settings. This involves first providing a starting prompt, which we generate in our experiments but can be passed in by a user, that we refer to as a seed prompt; then expanding on the seed prompt to generate increasingly more difficult and complex prompts. MT2IE-generated benchmarks are effective in that they match prior model rankings from benchmarks of handwritten prompts, while also being efficient in that they require far fewer examples than existing benchmarks. Additionally, MT2IE-generated benchmarks are dynamic, potentially addressing the dataset leakage and saturation issues of static benchmarks.
Iterative Prompt Generation.
Our approach iterates between two stages, shown on the right side of Figure figure 1. First, during iterative prompt generation, the MLLM is given an existing prompt $T_{i-1}$ and generates a new prompt $T_{i}$ that builds on $T_{i-1}$ to become more difficult. $T_{i}$ contains an additional object, spatial relationship, or object attribute compared to $T_{i-1}$, incrementally increasing the complexity of the prompt every iteration; see Figure figure 2 for an example. The seed prompt $T_{0}$ can be user-written or generated. The T2I model generates an image $I_{i}$ based on the prompt $T_{i}$. Next, during image scoring, the MLLM scores $I_{i}$ based on its faithfulness to prompt $T_{i}$. A full run consists of $N$ iterations of both iterative prompt generation and image scoring stages, with each iteration yielding a score $s_{i}$. We perform 756 full runs, with $N=5$ iterations per-run, resulting in a total of 3,780 scores from all runs. For these runs we evaluate a subset of 4 T2I models: CDM, LDM v2.1, LDM XL and LDM XL-Turbo. Our framework is flexible and supports any scoring metric during the image scoring stage; in this paper we use VQAScore (Lin et al., 2024), as it achieved the highest correlations with human judgment (see Experiment 1 in Section section 4.2).
Figure figure 3(a) shows average scores for each prompt iteration. We find that scores for all models decrease as prompt iterations increase, indicating that MT2IE effectively generates prompts that are progressively more challenging. To support this, we also analyze the change in linguistic properties of prompts across iterations in Figure figure 3(b); we use the Flesh-Kincaid score (Flesch, 1948) to compute semantic difficulty. As MT2IE-generated prompts get more difficult, they still maintain coherence and grammaticality. In general, prompts increase in their word count and semantic difficulty as prompt iterations occur. See an example progression of a 5-iteration run in Figure figure 2.
More Efficient Evaluation using our generated prompts.
After establishing that MT2IE generates progressively challenging prompt consistency benchmarks and achieves higher correlations in scoring generations, we next investigate how efficient MT2IE-generated benchmarks are compared to static benchmarks. Concretely, we ask: can we achieve the same evaluation results, i.e.,  same model rankings, while requiring fewer text prompts? We start with 4 generated seed prompts across 4 topic categories: household scenes (foods, household items, or furniture), descriptions of people, scenes with animals, and location descriptions. Each seed prompt is moderately complex (2-3 compositional concepts per prompt) and undergoes 4 difficulty progression iterations ($T_{0}\rightarrow T_{4}$), resulting in 20 total evaluation prompts in the generated benchmark (4 seeds $\times$ 5 prompts). Empirically we found that 5 prompts, resulting from 4 iterations, provided a sufficient amount of prompt difficulty (see Appendix appendix F for ablations on number of iterations). We refer to this as MT2IEbench20.
Table 3: Rank correlations between T2I model rankings on full GenAIBench (1600 prompts) versus only 20 prompts using our method versus all other evaluation methods (also see Figure figure 4). MT2IE produces notably high correlations rankings.
VQAScore CLIPscore VIEScore MT2IEi(Ours)
Kendall’s $\tau$ $-0.6428$ $0.3571$ $0.2857$ $0.8571$
Figure 4: Model rankings produced by MT2IE closely match GenAIBench’s ranking, while only using 20 generated prompts. Other methods produce mismatched model rankings when using 20 sampled GenAIBench prompts.
We rank the 8 T2I models listed in Section section 3 by (i) average score on 20 generated prompts from MT2IEbench20 and (ii) VQAScore evaluated on 1600 human-written prompts from GenAIBench (Li et al., 2024a). Next, for a direct comparison we randomly sample 20 prompts from GenAIBench, ensuring uniform sampling across all prompt categories, and refer to this as GenAIBench20. We rank the T2I models on GenAIBench20 by (iii) VQAScore (iv) CLIPScore (v) VIEScore.We run MT2IE-benchmark generation over 5 random seeds to ensure results are statistically significant (model score mean and standard deviations in Figure figure 7 and Figure figure 11, Appendix appendix C). Because variance across 5 seeds was extremely high for all metrics ran on GenAIBench20, all GenAIBench20 results are reported over 10 seeds. The relative rankings and scores for each evaluation method are shown in Figure figure 4.
We find that MT2IE, generating and scoring only 20 prompts per ranking, manages to closely match the T2I model ranking from the full, 1600-prompt GenAIBench benchmark. This indicates that MT2IE matches the T2I evaluation effectiveness of leading large-scale benchmarks while utilizing 20 times fewer prompts. Additionally, MT2IE-generated prompts eliminate the need for costly human-written prompt sets for benchmarking. All other evaluation methods tested failed to reliably match model rankings when applied to different random samples of 20 prompts from GenAIBench. For instance, VQAScore’s ranking on GenAIBench20 shows LDM v2.1 outperforming Flux, which is inconsistent with every other benchmark (Labs, 2024). These results show that simply randomly sampling a subset of prompts from existing datasets is not enough; our iterative approach generates strong examples that are able to truly probe performance.
6 Experiment 3: MLLMs as Adaptive Evaluator Agents
iteration 1: a set of knives and forks leaning against each other on a table Score: 0.912 Refer to caption iteration 2: a set of knives and forks leaning against each other on a table with a white tablecloth and a vase of flowers in the background
Score: 0.939
Refer to caption iteration 3: a set of knives and forks leaning against each other on a table with a white tablecloth and a vase of flowers in the background with a small wooden bowl of fruit on the table and a cat sitting on the floor next to the table
Score: 0.835
Refer to caption iteration 4: a set of knives and forks leaning against each other on a table with a white tablecloth and a vase of flowers in the background with a small wooden bowl of fruit on the table and a cat sitting on the floor next to the table and a bird perched on the tablecloth
Score: 0.386
Refer to caption iteration 5: a set of knives and forks leaning against each other on a table with a white tablecloth and a vase of flowers in the background [with a small wooden bowl of fruit on the table and a cat sitting on the floor next to the table and a bird perched on the tablecloth]
Score: 0.911
Figure 5: Example run with MT2IE’s Adaptive Evaluation, with generated images and alignment scores to illustrate the evaluation process. Portions in purple were added to the prompt from the previous iteration; portions in red were removed. Generations have high scores for iterations 1-3, so the prompts get progressively more difficult. When the score drops significantly at iteration 4, MT2IE reacts and simplifies the prompt.
We demonstrated that the MT2IE framework can iteratively generate prompts of increasing complexity, providing more efficient benchmarks compared to traditional static prompt sets. In this section, we move a step further in benchmark generation and ask Can MT2IE customize evaluations for individual T2I models based on their performance history? Individualized evaluation provides a nuanced understanding of differences between models and can facilitate the automatic discovery of model failure points. We explore adding an adaptive aspect to MT2IE, where the MLLM interacts with the T2I model by adjusting prompts based on the generated images seen from previous iterations.
Adaptive Prompt Generation.
(a) Prompt syntactic complexity, averaged across all prompts at each iteration. While most models get progressively more complex prompts from the first iteration, weaker models like LDM 2.1 struggle. All models hit a performance wall at the 4th iteration.
(b) Using MT2IE, we show how the scores for 8 T2I models change in our adaptive framework.
Figure 6: Results for MT2IE, where the MLLM iteratively generates prompts with feedback on individual model scores over 5 iterations. The difficulty of the generated prompt is dependent on the score from the previous iteration; description in Section 6.
As previously described, MT2IE iteratively expands upon prompts and scores corresponding image generations. In this adaptive setting, the MLLM additionally incorporates feedback about how well the T2I model performed on the previous iteration, via the score (Figure 1, right). This approach enables the MLLM to adaptively generate more challenging prompts when the T2I model performs well and simpler prompts when the model struggles. This dynamic adjustment is desirable because it enables discovery of failure patterns in strong models while avoiding uninformative results from excessively difficult prompts for weak models. By adjusting prompt difficulty dynamically, MT2IE addresses each scenario effectively.
In this setting, the MLLM is instructed to update prompt $T_{i}$ based on the prompt consistency score of the previous prompt $T_{i-1}$ according to:
  • score $0.0-0.2$: complexity halved
  • score $0.2-0.4$: complexity reduced by 1-2 terms
  • score $0.4-0.6$: prompt rephrased, same complexity
  • score $0.6-0.8$: complexity increased by 1-2 terms
  • score $0.8-1.0$: complexity increased by 2+ terms
With this system, T2I models that perform very poorly (score $\leq 0.2$) on prompt $T_{i}$ will have a significantly simpler next prompt $T_{i+1}$. Conversely, T2I models that are performing very well (score $\geq 0.8$) will have a significantly more difficult next prompt. To compute the final score after all $N$ iterations, each prompt’s score is weighted by its difficulty, ensuring greater rewards for models excelling at harder prompts compared to those scoring highly on easier ones. We show how the prompt complexity shifts across iterations for each model in Figure 6(a). Older, less advanced T2I models, e.g., LDM 2.1, show reduced prompt complexity almost immediately. Recent, high-performing T2I models, e.g., Flux, receive the most difficult prompts with MT2IE. An example of an adaptive run for CDM (Raffel et al., 2020) is shown in Figure 5.
Adaptive Prompting Preserves T2I Model Rankings.
Following our approach in Section 5, we use MT2IE to rank the same 8 T2I models on 20 adaptively generated prompts. We compare these rankings to GenAIBench’s rankings obtained from both its complete 1,600-prompt benchmark and a sampled 20-prompt subset, calculated using VQAScore from the CLIP-FlanT5 model. Because the adaptively generated prompts are custom to each model and therefore not directly comparable, we first standardize the scores to reflect prompt difficulty and then rank models with standardized scores. We use synactatic complexity to approximate prompt difficulty, computed using Yngve score which is a measure of the depth of a syntactic parse tree (Yngve, 1960). We present an exploration of various metrics to approximate prompt difficulty, including Flesch–Kincaid grade level (Flesch, 1948) and perplexity (Jelinek et al., 1977) in Appendix D. For each prompt-image pair, the initial consistency score is multiplied the prompt difficulty score to obtain the final score.
Figure 7: Our MT2IE framework produces T2I model rankings consistent with VQAScore on GenAIBench’s 1600 hand-written prompt benchmark (blue) with just 20 generated prompts in both the progressive difficulty setting (pink) and the model adaptive setting (green). When we downsample GenAIBench to 20 prompts (red), the rankings are noisy and uninformative, showing the effectiveness of our framework while using 80$\times$ less evaluation examples. Average rank (1=highest, 8=lowest) and rank variance over all seeds are plotted.
We show the average rank of each T2I model over all run seeds in Figure 7. Model rankings produced by MT2IE with adaptive prompting match the ranking from GenAIBench’s set of 1600 handwritten prompts, even when each individual model is evaluated on its bespoke prompt set. We again demonstrate the efficiency of our approach, requiring only 20 T2I generations to match evaluation results from an 80$\times$ larger, handwritten prompt benchmark.
7 Conclusion
Our work leverages the power of multimodal LLMs (MLLMs) as evaluator agents for text-to-image (T2I) model evaluation. Through MT2IE, we show that MLLMs are able to generate dynamic benchmarks of prompts. These benchmarks are effective for model evaluation, matching existing human-annotated rankings, and also highly efficient by requiring only a fraction of the examples. We also demonstrate that incorporating model-specific feedback to craft prompts enables the assessment of individual model’s abilities. Our MT2IE framework can be easily used with any off-the-shelf MLLM and incurs no computational cost on top of regular inference; MLLM system prompts and details on ease of adoption can be found in Appendices E, C, B.
Robust metrics and benchmarks are crucial to understanding and advancing T2I models. The current paradigm of T2I model evaluation relies heavily on legacy, static benchmarks (e.g., COCO (Lin et al., 2014), PartiPrompts (Yu et al., 2022)), which can be costly to create and quickly saturate (Ott et al., 2022; Gupta et al., 2024). MT2IE’s effectiveness shows that agentic behavior can be used in T2I evaluation to generate robust, efficient prompts to benchmark generation quality. We hope the community will build on these results to further develop dynamic and interactive evaluation frameworks.
Appendix A MT2IE For Aesthetic Quality Evaluations
To demonstrate MT2IE’s ability to generalize beyond evaluating image-prompt consistency, we use it to evaluate image aesthetic quality and compare its aesthetic quality judgments to a state-of-the-art aesthetics scoring model. MT2IE is easily adaptable to be used to evaluate other aspects of generated images by simply using a different set of MLLM system prompts; for aesthetic quality evaluation we use the system prompts in Figure figure 8. All other aspects of MT2IE are kept the same, and we use the 800 generated prompts and images from Section section 5. We compare aesthetic scores produced by MT2IE and the aesthetics predictor of LAION (2022), a popular aesthetics-scoring model (Kirstain et al., 2023; Wu et al., 2023b; Xu et al., 2023). MT2IE ’s ranking of 8 T2I models by average aesthetic score closely matches the pre-trained aesthetics-scoring model’s ranking, with a Kendall’s $\tau$ of $0.691$. Per-image aesthetic scores of 800 images are also positively correlated with a Kendall’s $\tau$ of $0.538$. These results, in tandem with our main paper experiments, show that MT2IE can effectively evaluate aesthetic quality and prompt consistency. Additionally, we demonstrate that MT2IE can be easily repurposed to evaluate axes other than prompt consistency by rewriting the MLLM prompt used for judgement.
You are a professional image judge. You will give a score between 0 and 10 to the given images about how visually pleasing they are. Here are things to consider when scoring the images:
How clear and sharp is the image?
Is anything in the image malformed?
Are the colors and figures in the image pleasing and coherent?
A score of 0 indicates that the image is very blurry, only has malformed objects, and unpleasant colors.
A score of 10 indicates that the image is clear, all objects are nicely depicted, and the colors are pleasant.
Score this image between 0 and 10. Only state the score and nothing else. (image)
Figure 8: Aesthetic quality evaluation system prompt (top) and user prompt given per-image evaluation (bottom).
Appendix B Ease of Adoption
The computational cost of one evaluation run with MT2IE is 40 inference passes of the utilized MLLM: 20 prompt generations with one prompt-image judgment each. Exact time and memory costs depend on MLLM size. However MT2IE greatly reduces the computational cost of generating images, as it is capable of matching T2I model rankings when evaluating on 20x less image-text pairs, which offsets the MLLM computation costs. MLLM system prompts for MT2IEare provided in Appendix appendix E and appendix C, and can be easily used with any off-the-shelf MLLM. Our code should be released upon acceptance.
Appendix C VQA Accuracy Scoring Results and MLLM System Prompts
In Section section 4, we validate that MLLMs are able to effectively score T2I models. We have two scoring paradigms: VQA Accuracy (question generation and answering) and VQAScore. VQA Accuracy requires generating multiple choice questions given a description. In the main paper, we only discuss VQAScore prompt alignment scoring as it acheived the highest correlation with human judgment in experiments in Section section 4. We therefore provide more details on the VQA Accuracy process below.
We provide the prompt given to each MLLM for question generation and validation (we only use questions for which the answer to the validation prompt is ”yes”). Figure 12 shows Llama-3.2-11B-Vision Instruction, Figure 13 shows Molmo-7b-d and Figure 14 shows Llava-v1.6-34b. We also provide experimental results for MT2IEwith both prompt alignment scoring paradigms discussed in Figures figure 10 and figure 11.
MT2IEwith both VQA Accuracy and VQAScore scoring are able to match GenAIBench’s T2I model rankings while using 80x less prompts. For MT2IEwith progressively harder prompts, T2I model scores from both VQA Accuracy and VQAScore scoring are extremely consistent if not exactly matching, showing that MLLM’s question generation and VQA abilities are on-par with the SOTA VQAScore metric.
Figure 9: All MLLMs obtain higher VQA accuracy than VQA models used as image judges in previous work. Results are on the 2k COCO image-caption pairs in TIFA160, as done in Hu et al. (2023).
Figure 10: Average rank of 8 T2I models (shown on x axis where 1=highest scoring, 8=lowest scoring) evaluated by MT2IEwith adaptive prompting and progressively difficult prompt generation, where each prompt generation method is run with both question generation and answering (VQA Accuracy) or VQAScore for prompt alignment scoring. All MT2IEsettings evaluate T2I models over 20 generated prompts, and we compare to ranks from GenAIBench with CLIP-FlanT5 VQAScore. MT2IErankings generally match GenAIBench’s T2I model rankings, while the Downsampled GenAIBench model rankings are uninformative as all models have roughly the same mean rank over all seeds and high variation. See Figure figure 4 for details on GenAIBench and Downsampled GenAIBench.
Figure 11: We evaluate 8 T2I models, comparing MT2IEwith progressively difficult prompt generation, and scoring prompt alignment with both question generation and answering (VQA Accuracy) or VQAScore, to other evaluation approaches. Scores from MT2IE are obtained from just 20 prompts that are both generated and scored by an MLLM. Scores from GenAIBench are obtained by generating and scoring 1600 human-written prompts. We find consistent T2I model rankings and scores between MT2IE, using only 20 prompts, and the full GenAIBench with both VQA Accuracy and VQAScore prompt alignment scoring methods (discussed in section 4). Just downsampling GenAIBench to 20 prompts breaks the relative model ranking.
Given an image description, generate multiple-choice questions that check if an image depicts everything in the image description.
You will not be given an image, just generate multiple-choice questions for the given image description and write nothing else.
First list the objects, attributes, spatial relations, and actions in the image description. Then write one questions for each of the stated items.
Write one question checking if each object is in the image. For any object attributes, write a question checking if the object in the image has the attribute.
For any spatial relations or actions, write a question checking if the spatial relation or action is shown in the image.
For each multiple-choice question, list answer choices and state the correct answer.
First write "Q:" and then state the question and nothing else.
Then on the next line write "Choices: " and then list the answer choices all on one line.
Then on the next line write "A: " and then state which answer out of the choices is correct.
All answer choices should be relevant to the question and short.
The only information you have is this image description: (prompt) Can this question be answered with only the image description? Question: (question) Answer yes or no and state nothing else.
Figure 12: For Llama-3.2-11B-Vision-Instruct, we provide the prompt for the question generation (top) and question validation (bottom).
Given an image description, generate multiple-choice questions that check if an image matches the image description.
You will not be given an image, just generate multiple-choice questions for the given image description and write nothing else.
For each multiple-choice question, generate answer choices and the correct answer.
Write one question for each object, attribute, and spatial relation in the image description.
Do not write questions about attributes of objects not mentioned in the description. Do not list more than 4 answer choices.
Generate questions, answer choices, and answers formatted like these examples:
Image description: a drawing of a red crab
Q: Is there a crab in the image?
Choices: yes, no
A: yes
Q: What animal is this?
Choices: lobster, fish, crab, eel
A: crab
Q: Is this a drawing?
Choices: yes, no
A: yes
Q: What color is the crab?
Choices: grey, blue, red, purple
A: red
Image description: A cheese grater with a red handle sitting on a white countertop next to a yellow cutting board and a green apple
Q: Is there a cheese grater?
Choices: yes, no
A: yes
Q: What kitchen tool is on the counter?
Choices: blender, coffee maker, cheese grater, mixer
A: cheese grater
Q: What color is the handle of the cheese grater?
Choices: green, silver, white, red
A: red
Q: What color is the countertop?
Choices: black, white, marble, grey
A: white
Q: Is there a cutting board?
Choices: yes, no
A: yes
Q: What color is the cutting board?
Choices: white, yellow, blue, brown
A: yellow
Q: Is there an apple?
Choices: yes, no
A: yes
Q: What color is the apple?
Choices: yellow, red, green
A: green
Image description: laundry detergent
Q: Is there laundry detergent in the image?
Choices: yes, no
A: yes
Image description: a tropical rainforest with a waterfall cascading down a cliff a group of monkeys swinging from the trees and a large colorful parrot perched on a branch
Q: Is there a tropical rainforest?
Choices: yes, no
A: yes
Q: Is there a waterfall cascading down a cliff?
Choices: yes, no
A: yes
Q: Are there a group of monkeys in the image?
Choices: yes, no
A: yes
Q: What are the monkeys doing?
Choices: eating, swinging from trees, swimming, sleeping
A: swinging from trees
Q: Is there a parrot in the image?
Choices: yes, no
A: yes
Q: Where is the parrot in the image?
Choices: flying overhead, perched on a branch, standing on a rock, flying above the water
A: perched on a branch
Given this image caption: (prompt)
Does this question ask about information explicitly included in the caption: (question)
Answer yes or no and state nothing else.
Figure 13: For Molmo-7b-d, we provide the prompt for the question generation (top box) and question validation (bottom box).
Given an image description, generate multiple-choice questions that check if an image matches the image description.
You will not be given an image, just generate multiple-choice questions for the given image description and write nothing else.
For each multiple-choice question, generate answer choices and the correct answer.
Write one question for each object, attribute, and spatial relation in the image description.
Do not write questions about attributes of objects not mentioned in the description. Do not list more than 4 answer choices.
Generate questions, answer choices, and answers formatted like these examples:
Image description: a drawing of a red crab
Q: Is there a crab in the image?
Choices: yes, no
A: yes
Q: What animal is this?
Choices: lobster, fish, crab, eel
A: crab
Q: Is this a drawing?
Choices: yes, no
A: yes
Q: What color is the crab?
Choices: grey, blue, red, purple
A: red
Image description: A cheese grater with a red handle sitting on a white countertop next to a yellow cutting board and a green apple
Q: Is there a cheese grater?
Choices: yes, no
A: yes
Q: What kitchen tool is on the counter?
Choices: blender, coffee maker, cheese grater, mixer
A: cheese grater
Q: What color is the handle of the cheese grater?
Choices: green, silver, white, red
A: red
Q: What color is the countertop?
Choices: black, white, marble, grey
A: white
Q: Is there a cutting board?
Choices: yes, no
A: yes
Q: What color is the cutting board?
Choices: white, yellow, blue, brown
A: yellow
Q: Is there an apple?
Choices: yes, no
A: yes
Q: What color is the apple?
Choices: yellow, red, green
A: green
Image description: laundry detergent
Q: Is there laundry detergent in the image?
Choices: yes, no
A: yes
Figure 14: For Llava-v1.6-34b, we provide the prompt for the question generation (top) and question validation (bottom).
Appendix D Analysis of Prompt Difficulty Metrics
For MT2IEwith Adaptive Prompting, discussed in Section section 6, we multiply prompt alignment scores with a prompt difficulty metric, in order to standardize prompt alignment scores since all models are no longer evaluated on the same prompts. An ideal prompt difficulty metric reflects how difficult it is for a T2I model to generate an image faithful to the prompt, and can weight the final prompt alignment score appropriately: scaling difficult prompts to higher values and easy prompts to lower values.
However to our knowledge, there is no existing work exploring metrics for T2I prompt difficulty. So we considered various NLP metrics for language complexity as prompt difficulty proxy metrics, including: word count, syllable count, Yngve score (Yngve, 1960) (a measure of the depth of a syntactic parse tree), Flesch–Kincaid scores (Flesch, 1948), and perplexity (Jelinek et al., 1977). We did preliminary analysis on a small set of MLLM-generated prompts and VQAScores and found that Yngve score was the metric most correlated with VQAScore, so we used Yngve score as the prompt difficulty weight in Section section 6’s experiments. After conducting all our experiments, we now have 20,640 (MLLM-generated) prompts along with their corresponding T2I model VQAScores, to analyze in order to identify a potential metric for prompt difficulty.
Table 4: Prompt metric correlations of 20,640 T2I model VQAScores of prompts and their generated images. The higher the correlation, the more viable the prompt metric is as a prompt difficulty metrics. Perplexity was calculated with Llama2, and we pre-pended ”There’s an image of a” to the prompt as perplexity values were extremely high (in the thousands) for many prompts without the prefix.
Prompt Metric Kendall’s $\tau$ Spearman’s $\rho$
Flesch-Kincaid Grad Level $-0.1368$ $-0.1011$
Word Count $-0.1511$ $-0.1188$
Syllable Count $-0.1437$ $-0.1076$
Average Syllables per Word $0.0476$ $0.1022$
Average Word Length $0.1025$ $0.1739$
Perplexity (Llama2) $0.0574$ $0.0407$
Yngve Score $-0.0970$ $-0.0965$
Table table 4 shows correlations between various metrics calculated on prompts and T2I model VQAScores for those same prompts. Interestingly, simple statistics such as average word length and syllable count perform the best, outperforming semantic complexity metrics like Flesh-Kincaid grade level or Yngve score. Effectively measuring prompt difficulty is an interesting area of future work that we believe is very useful to the field of T2I evaluation.
Appendix E MLLM System Prompts for Prompt Generation
MT2IE Iterative Difficulty Prompt Generation (Section section 5)
For prompt generation in Section section 5, we use the system prompt shown in Figure figure 15 and run generation with temperature 0.3 and no top-k sampling.
You are responsible for rewriting prompts for a computer program that generates images.
You will be given an existing prompt that you must add complexity to by adding one term that is an additional object, a new spatial relationship, or new attribute.
Do not add subjective descriptors to the prompt, it should be as factual as possible. Do not use fluffy, poetic language, or any words beyond the elementary school level.
Every time you will be givng an existing prompt, write "Prompt:" and write the new prompt, do not state anything else.
Existing prompt: "a serene mountain lake with a small wooden dock and a small wooden cabin on the shore"
Prompt: a serene mountain lake with a small wooden dock and a small wooden cabin on the shore, with birds in the sky
Existing prompt: "elderly man wearing a backpack and a hat walking with a cane and a dog on a leash"
Prompt: elderly man wearing a backpack and a hat walking with a cane and a big white dog on a leash
Existing prompt: "a green fish hiding in a coral reef next to a crab"
Prompt: a green fish hiding in a coral reef next to a crab and a purple octopus
Existing prompt: (previous prompt)
Figure 15: System prompt for Llava-v1.6-34b when generating iteratively more difficult prompts.
MT2IE Adaptive Prompting Prompt Generation (Section section 6)
For adaptive prompt generation in Section section 6, we use the system prompts shown in Figures figure 16, figure 17, figure 18 and run generation with temperature 0.3 and no top-k sampling.
Score $[0.0,0.2]$:
You are responsible for rewriting prompts for a computer program that generates images.
You will be given an existing prompt that you must simplify so that there are half the amount of terms.
Remove objects, attributes, or spatial relationships in the existing prompt and rewrite the prompt so that the new prompt makes sense but only describes half the amount of concepts.
If the existing prompt is too simple to reduce, just rewrite the prompt to be clearer.
Every time you will be givng an existing prompt, write "Prompt:" and write the new prompt, do not state anything else.
Existing prompt: "a serene mountain lake with a small wooden dock a few birds flying overhead and a small wooden cabin on the shore"
Prompt: a mountain lake with a wooden dock
Existing prompt: "elderly man walking with a cane a dog on a leash a backpack on his back a hat on his head and a pair of sunglasses on his face" Prompt: an elderly man walking with a cane and a dog
Existing prompt: "green fish swimming in a coral reef next to a purple octopus" Prompt: a green fish in a coral reef
Score $(0.2,0.4]$:
You are responsible for rewriting prompts for a computer program that generates images.
You will be given an existing prompt that you must simplify by removing two terms.
The terms to remove are objects, attributes, or spatial relationships in the existing prompt.
If there are less than two terms in the existing prompt, just rewrite the prompt to be simpler.
Every time you will be givng an existing prompt, write "Prompt:" and write the new prompt, do not state anything else.
Existing prompt: "a serene mountain lake with a small wooden dock a few birds flying overhead and a small wooden cabin on the shore"
Prompt: a mountain lake with a small wooden dock and a few birds
Existing prompt: "elderly man walking with a cane a dog on a leash a backpack on his back a hat on his head and a pair of sunglasses on his face"
Prompt: elderly man with a cane wearing a backpack and a dog on a leash
Existing prompt: "a green, striped fish hiding in a coral reef next to a crab"
Prompt: a green fish in a coral reef
Figure 16: System prompts for Llava-v1.6-34b when generating adaptive prompts, for each of the score bins listed.
Score $(0.4,0.6)$:
You are responsible for rewriting prompts for a computer program that generates images.
You will be given an existing prompt that you must rewrite to make more clear and simple.
Rewrite the prompt with simpler words, clearer phrasing, and add punctuation if needed. The meaning of the new and existing prompts must be the same.
Every time you will be givng an existing prompt, write "Prompt:" and write the new prompt, do not state anything else.
Existing prompt: "a serene mountain lake with a small wooden dock a few birds flying overhead and a small wooden cabin on the shore"
Prompt: a lake on a mountain with a small dock, birds flying in the sky, and a small cabin on the shore
Existing prompt: "elderly man walking with a cane a dog on a leash a backpack on his back a hat on his head and a pair of sunglasses on his face"
Prompt: an old man wearing a backpack, a hat, and sunglasses walking with a cane and his dog on a leash
Existing prompt: "a group of seals jumping out of the water in a large body of water with a boat nearby and a person on the boat taking a picture while a seagull is flying overhead"
Prompt: seals jumping out of a large body of water, near a boat with a person taking a picture onboard and a seagull flying overhead
Existing prompt: "a city with crowded streets and skyscrapers with a mountain in the background and traffic in the foreground and a large stadium in the distance"
Prompt: a city with busy streets and tall buildings and a mountain in the background and cars in the foreground, with a large stadium in the distance
Score $[0.6,0.8)$:
You are responsible for rewriting prompts for a computer program that generates images.
You will be given an existing prompt that you must add complexity to by adding one term that is an additional object, a new spatial relationship, or new attribute.
Do not add subjective descriptors to the prompt, it should be as factual as possible. Do not use fluffy, poetic language, or any words beyond the elementary school level.
Every time you will be givng an existing prompt, write "Prompt:" and write the new prompt, do not state anything else.
Existing prompt: "a serene mountain lake with a small wooden dock and a small wooden cabin on the shore"
Prompt: a serene mountain lake with a small wooden dock and a small wooden cabin on the shore, with birds in the sky
Existing prompt: "elderly man wearing a backpack and a hat walking with a cane and a dog on a leash"
Prompt: elderly man wearing a backpack and a hat walking with a cane and a big white dog on a leash
Existing prompt: "a green fish hiding in a coral reef next to a crab"
Prompt: a green fish hiding in a coral reef next to a crab and a purple octopus
Figure 17: System prompts for Llava-v1.6-34b when generating adaptive prompts, for each of the score bins listed.
Score $[0.8,1.0]$:
You are responsible for rewriting prompts for a computer program that generates images.
You will be given an existing prompt that you must add complexity to by adding a few terms. Each term is an additional object, a new spatial relationship, or new attribute.
Do not add subjective descriptors to the prompt, it should be as factual as possible. Do not use fluffy, poetic language, or any words beyond the elementary school level.
Every time you will be givng an existing prompt, write "Prompt:" and write the new prompt, do not state anything else.
Existing prompt: "a serene mountain lake with a small wooden dock and a cabin on the shore"
Prompt: a serene mountain lake with a small wooden dock and a large cabin on the shore, with birds in the sky and a fisherman fishing on the dock
Existing prompt: "elderly man wearing a backpack and a hat walking with a cane and a dog on a leash"
Prompt: elderly man wearing a blue backpack, a red hat, and sunglasses walking with a cane and a big white dog on a leash
Existing prompt: "a green fish hiding in a coral reef next to a crab"
Prompt: a striped green fish hiding in a coral reef next to a crab and a purple octopus with the open ocean in the background
Existing prompt: "a lemon tree on a hill"
Prompt: a big lemon tree on a hill with a dog sitting underneath
Figure 18: System prompts for Llava-v1.6-34b when generating adaptive prompts, for the score bin listed.
Appendix F Ablations on Prompt Iterations
Table 5: Kendall’s $\tau$ rank correlations between T2I model rankings produced by MT2IEfor all prompt iteration values and the full, 1600 prompt GenAIBench benchmark. We show the average $\tau$ between MT2IEwith Iterative Difficulty and Adaptive Prompting for each prompt iteration value, and compare to the ranking produced by downsampling GenAIBench to the same number of prompts as our highest-correlated prompt iteration value. MT2IEwith all prompt iteration values produces rankings notably more correlated to GenAIBench than downsampled GenAIBench.
T2I Evaluation Method Kendall’s $\tau$
Downsampled GenAIBench (20 Prompts) $-0.8693$
MT2IE: 2 Prompt Iterations (8 Prompts) $0.6672$
MT2IE: 5 Prompt Iterations (20 Prompts) $0.7275$
MT2IE: 10 Prompt Iterations (40 Prompts) $0.7004$
MT2IE: 12 Prompt Iterations (48 Prompts) $0.7025$
MT2IEgenerates a configurable number of prompts: in the Iterative Difficulty setting discussed in Section section 5 pronpts are progressively more difficulty in each iteration, whereas in the Adaptive Prompting setting discussed in Section section 6 each iteration’s prompt is rewritten by the MLLM based on the previous prompt’s alignment score.
Rankings of 8 T2I models for prompt iterations={2, 5, 10, 12} for Iterative Difficulty are shown in Figure figure 19 and Adaptive Prompting in Figure figure 20. Additionally we calculate Kendall’s $\tau$ rank correlation coefficient for all T2I model rankings, using the ranking obtained from GenAIBench as the ground truth ranking, shown in Table table 5. We found that using 5 prompt iterations, i.e.,  20 total prompts, with MT2IEyields strongly-correlated model scores with the GenAIBench benchmark which contains 1600 prompts. 5 prompt iterations results in the highest ranking correlation and smaller standard deviations in model rankings. We suspect that using more prompt iterations in the Iterative Difficulty setting can result in prompts that are too hard for many T2I models, resulting in uninformative scores and model rankings.
Figure 19: Average rank of 8 T2I models (shown on x axis where 1=highest scoring, 8=lowest scoring) evaluated by MT2IEwith progressively difficult prompt generation with various prompt iterations. All prompt iteration counts yield more informative rankings that better match GenAIBench than downsampled GenAIBench.
Figure 20: Average rank of 8 T2I models (shown on x axis where 1=highest scoring, 8=lowest scoring) evaluated by MT2IEwith adaptive prompt generation with various prompt iterations. All prompt iteration counts yield more informative rankings that better match GenAIBench than downsampled GenAIBench.
Appendix G Comparing TIFAv1.0 and the Llava-Generated Question Set from TIFAv1.0 Prompts
Figure 21: Element mismatch counts for all TIFAv1.0 4,081 captions (prompts). Most captions have just 1 or no element mismatch, indicating that Llava can break down and generate questions for prompts with similar performance to the proprietary GPT-3.
We analyze and compare TIFAv1.0 and our Llava-Generated question set (Llava-Gen TIFA) for the same 4,081 prompts from TIFA. This is done to verify that Llava is capable of generating questions as effectively as proprietary LLMs (TIFAv1.0 uses GPT-3 to generate questions).
General question set statistics
Overall, TIFAv1.0 has an average of 6.87 questions per prompt and Llava-Gen TIFA has an average of 6.14 questions per prompt. This results in Llava-Gen TIFA containing around 10% less questions than TIFAv1.0 while covering the same number of prompts. Despite this, Llava-Gen TIFA achieves higher correlations with human judgment than TIFAv1.0, supporting our claims that MLLM-generated T2I benchmarks are more efficient. The distribution of question counts per prompt are shown in Figure figure 21.
Similarity of question elements
TIFAv1.0 provides entity tags for each of their prompts, indicating which prompt elements the generated questions cover (Hu et al., 2023). We prompt Llava to similarly identify elements when generating questions and analyze the elements captured in TIFAv1.0 vs. Llava-Gen TIFA.
Figure 22: Element mismatch counts for all TIFAv1.0 4,081 captions (prompts). Most captions have just 1 or no element mismatch, indicating that Llava can break down and generate questions for prompts with similar performance to the proprietary GPT-3.
Across all prompts, TIFAv1.0 has an average of 4.68 elements per prompt and Llava-Gen TIFA has 4.13 elements per prompt, indicating an extremely similar number of elements identified for each prompt. For prompts with mismatched elements, the average number of mismatched elements is only 1.13, and we show counts of prompts with mismatched elements in Figure figure 22. The average BLEU (Papineni et al., 2002) score between TIFAv1.0 and Llava-Gen TIFA questions about the same elements is 0.65, which indicates a high level of similarity between questions covering the same concepts.

### v2

@@
 We use this MLLM-based approach to compare the prompt-generation consistency of 8 state-of-the-art T2I models and find that the rankings from our approach are comparable to those reported by static benchmarks, while only using a fraction ($\sim$1%) of prompts.
-We also show that MT2IE is easily generalizable to evaluate aspects of image generation beyond prompt-generation consistency, with results for aesthetic quality presented in Appendix appendix A.
+We also show that MT2IE is easily generalizable to evaluate aspects of image generation beyond prompt-generation consistency, with results for aesthetic quality presented in Appendix A.
 Finally, we investigate MLLMs as evaluator agents for prompt-generation consistency by enabling the interaction between the MLLM and the T2I model.
@@
 In the following sections, we focus on evaluating image-text consistency as it directly impacts the overall visual quality of generated images (Astolfi et al., 2024; Xu et al., 2023; Kirstain et al., 2023).
-Additionally, in Appendix appendix A we show that MT2IE is an easily generalizable framework, capable of evaluating other aspects of T2I models such as image aesthetics.
+Additionally, in Appendix A we show that MT2IE is an easily generalizable framework, capable of evaluating other aspects of T2I models such as image aesthetics.
 We use MT2IE to evaluate a total of 8 recent T2I models including latent diffusion model (LDM) v2.1 (Rombach et al., 2022), LDM v3 (Esser et al., 2024), LDM XL (Podell et al., 2023), LDM XL-Turbo (Sauer et al., 2025), Flux (Labs, 2024), a cascaded pixel-based diffusion model (CDM) (Raffel et al., 2020), Playground 2.5 (Li et al., 2024b) and Pixart-$\alpha$ (Chen et al., 2023).
@@
 We evaluate using TIFA160 (Hu et al., 2023), and show that a single open-sourced MLLM can recreate this benchmark and score questions with better correlation to human judgment.
-We detail this experiment in Section section 4.
+We detail this experiment in Section 4.
 Experiment 2: MLLMs to generate prompt consistency benchmarks.
@@
 We also show that our method is efficient, accurately ranking models with significantly fewer prompts than existing benchmarks of handwritten prompts.
-We detail these experiments in Section section 5.
+We detail these experiments in Section 5.
 Experiment 3: MLLMs to generate interactive prompt consistency benchmarks.
@@
 Specifically, we show that MLLMs can adjust the complexity of prompts based on model performance on previous prompts.
-This enables each individual T2I model to have bespoke prompts tailored to its abilities, discussed in Section section 6.
+This enables each individual T2I model to have bespoke prompts tailored to its abilities, discussed in Section 6.
 4 Experiment 1: Validation on Static Benchmarks & Metrics
@@
 We compare the MLLMs’ VQA accuracy to the 2 best VQA models from Hu et al.
-(2023); results are in Figure figure 9 (Appendix appendix C) and show that all 3 MLLMs outperform the VQA models used in existing GQA benchmarks.
+(2023); results are in Figure 9 (Appendix C) and show that all 3 MLLMs outperform the VQA models used in existing GQA benchmarks.
 Table 1: Human judgment correlations of existing VQA-based T2I consistency benchmarks and MT2IE, measured by Spearman’s $\rho$ and Kendall’s $\tau$ on TIFA160.
@@
 We use all 3 MLLMs to generate and score questions about all prompts in TIFA160.
-The system prompts used to generate questions is shown in Appendix appendix C.
+The system prompts used to generate questions is shown in Appendix C.
 To assess quality of the generated questions, we follow previous work and measure the correlation between the VQA accuracy (from LLM-generated questions from existing benchmarks (Cho et al., 2024; Hu et al., 2023; Gupta and Kembhavi, 2022) or MLLM-generated questions in our work) and human ratings of image-text consistency using a 1-5 Likert scale.
 We use images and human ratings from prior work (Cho et al., 2024) on 5 different T2I models: LDM1.1 (Rombach et al., 2022), LDM1.5 (Rombach et al., 2022), LDM2.1 (Rombach et al., 2022), minDALL-E (Saehoon et al., 2021), and VQ-Diffusion (Gu et al., 2022).
-We show the Spearman’s $\rho$ and Kendall’s $\tau$ in Table table 1.
+We show the Spearman’s $\rho$ and Kendall’s $\tau$ in Table 1.
 Notably, every MLLM achieves higher correlation with human annotations than existing GQA-based approaches.
@@
 We analyze question statistics and BLEU score (Papineni et al., 2002) (a commonly used NLP metric to compare two sets of text) and find that entities in both sets of questions are very similar.
-Detailed results on generated question set verification are in Appendix appendix G.
+Detailed results on generated question set verification are in Appendix G.
 4.2 VQAScore Approach
@@
 We compare the correlations with human ratings on TIFA160 between VQAScore calculated using the 3 aforementioned MLLMs, VQAScore calculated with the CLIP-FlanT5 model (which was fine-tuned for SOTA VQAScore calculation (Lin et al., 2024)), and CLIPScore.
-We use the same human ratings as in Section section 4.1; results are in Table table 2.
+We use the same human ratings as in Section 4.1; results are in Table 2.
 All MLLMs outperform CLIPScore in human correlation, but only Llava-v1.6-34b achieves higher correlation with human ratings than CLIP-FlanT5.
@@
 Llava-v1.6-34b achieved the best correlation with human ratings when used to compute VQAScore as prompt consistency scores; so we use Llava-v1.6-34b with VQAScore as our MLLM evaluator for further experiments.
-Experiments with question generation and answering for prompt consistency scoring are in Appendix appendix C.
+Experiments with question generation and answering for prompt consistency scoring are in Appendix C.
 This validation lays the framework for our experiments, demonstrating that open-sourced MLLMs can be used for end-to-end T2I evaluation.
@@
 Figure 3: Results for MT2IE, where the MLLM iteratively generates progressively more difficult prompts over 5 iterations.
-In Section section 4, we showed that MLLMs are able to reproduce static benchmarks using both question-answering and prompt likelihood approaches.
+In Section 4, we showed that MLLMs are able to reproduce static benchmarks using both question-answering and prompt likelihood approaches.
 We now introduce our approach, the MT2IE framework, and show that it can be used to iteratively generate new, efficient benchmarks for image-text consistency.
@@
 Iterative Prompt Generation.
-Our approach iterates between two stages, shown on the right side of Figure figure 1.
+Our approach iterates between two stages, shown on the right side of Figure 1.
 First, during iterative prompt generation, the MLLM is given an existing prompt $T_{i-1}$ and generates a new prompt $T_{i}$ that builds on $T_{i-1}$ to become more difficult.
-$T_{i}$ contains an additional object, spatial relationship, or object attribute compared to $T_{i-1}$, incrementally increasing the complexity of the prompt every iteration; see Figure figure 2 for an example.
+$T_{i}$ contains an additional object, spatial relationship, or object attribute compared to $T_{i-1}$, incrementally increasing the complexity of the prompt every iteration; see Figure 2 for an example.
 The seed prompt $T_{0}$ can be user-written or generated.
@@
 For these runs we evaluate a subset of 4 T2I models: CDM, LDM v2.1, LDM XL and LDM XL-Turbo.
-Our framework is flexible and supports any scoring metric during the image scoring stage; in this paper we use VQAScore (Lin et al., 2024), as it achieved the highest correlations with human judgment (see Experiment 1 in Section section 4.2).
-Figure figure 3(a) shows average scores for each prompt iteration.
+Our framework is flexible and supports any scoring metric during the image scoring stage; in this paper we use VQAScore (Lin et al., 2024), as it achieved the highest correlations with human judgment (see Experiment 1 in Section 4.2).
+Figure 3(a) shows average scores for each prompt iteration.
 We find that scores for all models decrease as prompt iterations increase, indicating that MT2IE effectively generates prompts that are progressively more challenging.
-To support this, we also analyze the change in linguistic properties of prompts across iterations in Figure figure 3(b); we use the Flesh-Kincaid score (Flesch, 1948) to compute semantic difficulty.
+To support this, we also analyze the change in linguistic properties of prompts across iterations in Figure 3(b); we use the Flesh-Kincaid score (Flesch, 1948) to compute semantic difficulty.
 As MT2IE-generated prompts get more difficult, they still maintain coherence and grammaticality.
 In general, prompts increase in their word count and semantic difficulty as prompt iterations occur.
-See an example progression of a 5-iteration run in Figure figure 2.
+See an example progression of a 5-iteration run in Figure 2.
 More Efficient Evaluation using our generated prompts.
@@
 Each seed prompt is moderately complex (2-3 compositional concepts per prompt) and undergoes 4 difficulty progression iterations ($T_{0}\rightarrow T_{4}$), resulting in 20 total evaluation prompts in the generated benchmark (4 seeds $\times$ 5 prompts).
-Empirically we found that 5 prompts, resulting from 4 iterations, provided a sufficient amount of prompt difficulty (see Appendix appendix F for ablations on number of iterations).
+Empirically we found that 5 prompts, resulting from 4 iterations, provided a sufficient amount of prompt difficulty (see Appendix F for ablations on number of iterations).
 We refer to this as MT2IEbench20.
-Table 3: Rank correlations between T2I model rankings on full GenAIBench (1600 prompts) versus only 20 prompts using our method versus all other evaluation methods (also see Figure figure 4).
+Table 3: Rank correlations between T2I model rankings on full GenAIBench (1600 prompts) versus only 20 prompts using our method versus all other evaluation methods (also see Figure 4).
 MT2IE produces notably high correlations rankings.
@@
 Other methods produce mismatched model rankings when using 20 sampled GenAIBench prompts.
-We rank the 8 T2I models listed in Section section 3 by (i) average score on 20 generated prompts from MT2IEbench20 and (ii) VQAScore evaluated on 1600 human-written prompts from GenAIBench (Li et al., 2024a).
+We rank the 8 T2I models listed in Section 3 by (i) average score on 20 generated prompts from MT2IEbench20 and (ii) VQAScore evaluated on 1600 human-written prompts from GenAIBench (Li et al., 2024a).
 Next, for a direct comparison we randomly sample 20 prompts from GenAIBench, ensuring uniform sampling across all prompt categories, and refer to this as GenAIBench20.
-We rank the T2I models on GenAIBench20 by (iii) VQAScore (iv) CLIPScore (v) VIEScore.We run MT2IE-benchmark generation over 5 random seeds to ensure results are statistically significant (model score mean and standard deviations in Figure figure 7 and Figure figure 11, Appendix appendix C).
+We rank the T2I models on GenAIBench20 by (iii) VQAScore (iv) CLIPScore (v) VIEScore.We run MT2IE-benchmark generation over 5 random seeds to ensure results are statistically significant (model score mean and standard deviations in Figure 7 and Figure 11, Appendix C).
 Because variance across 5 seeds was extremely high for all metrics ran on GenAIBench20, all GenAIBench20 results are reported over 10 seeds.
-The relative rankings and scores for each evaluation method are shown in Figure figure 4.
+The relative rankings and scores for each evaluation method are shown in Figure 4.
 We find that MT2IE, generating and scoring only 20 prompts per ranking, manages to closely match the T2I model ranking from the full, 1600-prompt GenAIBench benchmark.
@@
 To demonstrate MT2IE’s ability to generalize beyond evaluating image-prompt consistency, we use it to evaluate image aesthetic quality and compare its aesthetic quality judgments to a state-of-the-art aesthetics scoring model.
-MT2IE is easily adaptable to be used to evaluate other aspects of generated images by simply using a different set of MLLM system prompts; for aesthetic quality evaluation we use the system prompts in Figure figure 8.
-All other aspects of MT2IE are kept the same, and we use the 800 generated prompts and images from Section section 5.
+MT2IE is easily adaptable to be used to evaluate other aspects of generated images by simply using a different set of MLLM system prompts; for aesthetic quality evaluation we use the system prompts in Figure 8.
+All other aspects of MT2IE are kept the same, and we use the 800 generated prompts and images from Section 5.
 We compare aesthetic scores produced by MT2IE and the aesthetics predictor of LAION (2022), a popular aesthetics-scoring model (Kirstain et al., 2023; Wu et al., 2023b; Xu et al., 2023).
@@
 However MT2IE greatly reduces the computational cost of generating images, as it is capable of matching T2I model rankings when evaluating on 20x less image-text pairs, which offsets the MLLM computation costs.
-MLLM system prompts for MT2IEare provided in Appendix appendix E and appendix C, and can be easily used with any off-the-shelf MLLM.
+MLLM system prompts for MT2IEare provided in Appendix E and C, and can be easily used with any off-the-shelf MLLM.
 Our code should be released upon acceptance.
 Appendix C VQA Accuracy Scoring Results and MLLM System Prompts
-In Section section 4, we validate that MLLMs are able to effectively score T2I models.
+In Section 4, we validate that MLLMs are able to effectively score T2I models.
 We have two scoring paradigms: VQA Accuracy (question generation and answering) and VQAScore.
 VQA Accuracy requires generating multiple choice questions given a description.
-In the main paper, we only discuss VQAScore prompt alignment scoring as it acheived the highest correlation with human judgment in experiments in Section section 4.
+In the main paper, we only discuss VQAScore prompt alignment scoring as it acheived the highest correlation with human judgment in experiments in Section 4.
 We therefore provide more details on the VQA Accuracy process below.
 We provide the prompt given to each MLLM for question generation and validation (we only use questions for which the answer to the validation prompt is ”yes”).
-Figure 12 shows Llama-3.2-11B-Vision Instruction, Figure 13 shows Molmo-7b-d and Figure 14 shows Llava-v1.6-34b.
-We also provide experimental results for MT2IEwith both prompt alignment scoring paradigms discussed in Figures figure 10 and figure 11.
+12 shows Llama-3.2-11B-Vision Instruction, 13 shows Molmo-7b-d and 14 shows Llava-v1.6-34b.
+We also provide experimental results for MT2IEwith both prompt alignment scoring paradigms discussed in Figures 10 and 11.
 MT2IEwith both VQA Accuracy and VQAScore scoring are able to match GenAIBench’s T2I model rankings while using 80x less prompts.
@@
 MT2IErankings generally match GenAIBench’s T2I model rankings, while the Downsampled GenAIBench model rankings are uninformative as all models have roughly the same mean rank over all seeds and high variation.
-See Figure figure 4 for details on GenAIBench and Downsampled GenAIBench.
+See Figure 4 for details on GenAIBench and Downsampled GenAIBench.
 Figure 11: We evaluate 8 T2I models, comparing MT2IEwith progressively difficult prompt generation, and scoring prompt alignment with both question generation and answering (VQA Accuracy) or VQAScore, to other evaluation approaches.
@@
 Scores from GenAIBench are obtained by generating and scoring 1600 human-written prompts.
-We find consistent T2I model rankings and scores between MT2IE, using only 20 prompts, and the full GenAIBench with both VQA Accuracy and VQAScore prompt alignment scoring methods (discussed in section 4).
+We find consistent T2I model rankings and scores between MT2IE, using only 20 prompts, and the full GenAIBench with both VQA Accuracy and VQAScore prompt alignment scoring methods (discussed in 4).
 Just downsampling GenAIBench to 20 prompts breaks the relative model ranking.
@@
 Appendix D Analysis of Prompt Difficulty Metrics
-For MT2IEwith Adaptive Prompting, discussed in Section section 6, we multiply prompt alignment scores with a prompt difficulty metric, in order to standardize prompt alignment scores since all models are no longer evaluated on the same prompts.
+For MT2IEwith Adaptive Prompting, discussed in Section 6, we multiply prompt alignment scores with a prompt difficulty metric, in order to standardize prompt alignment scores since all models are no longer evaluated on the same prompts.
 An ideal prompt difficulty metric reflects how difficult it is for a T2I model to generate an image faithful to the prompt, and can weight the final prompt alignment score appropriately: scaling difficult prompts to higher values and easy prompts to lower values.
@@
 So we considered various NLP metrics for language complexity as prompt difficulty proxy metrics, including: word count, syllable count, Yngve score (Yngve, 1960) (a measure of the depth of a syntactic parse tree), Flesch–Kincaid scores (Flesch, 1948), and perplexity (Jelinek et al., 1977).
-We did preliminary analysis on a small set of MLLM-generated prompts and VQAScores and found that Yngve score was the metric most correlated with VQAScore, so we used Yngve score as the prompt difficulty weight in Section section 6’s experiments.
+We did preliminary analysis on a small set of MLLM-generated prompts and VQAScores and found that Yngve score was the metric most correlated with VQAScore, so we used Yngve score as the prompt difficulty weight in Section 6’s experiments.
 After conducting all our experiments, we now have 20,640 (MLLM-generated) prompts along with their corresponding T2I model VQAScores, to analyze in order to identify a potential metric for prompt difficulty.
@@
 Yngve Score $-0.0970$ $-0.0965$
-Table table 4 shows correlations between various metrics calculated on prompts and T2I model VQAScores for those same prompts.
+Table 4 shows correlations between various metrics calculated on prompts and T2I model VQAScores for those same prompts.
 Interestingly, simple statistics such as average word length and syllable count perform the best, outperforming semantic complexity metrics like Flesh-Kincaid grade level or Yngve score.
@@
 Appendix E MLLM System Prompts for Prompt Generation
-MT2IE Iterative Difficulty Prompt Generation (Section section 5)
-For prompt generation in Section section 5, we use the system prompt shown in Figure figure 15 and run generation with temperature 0.3 and no top-k sampling.
+MT2IE Iterative Difficulty Prompt Generation (Section 5)
+For prompt generation in Section 5, we use the system prompt shown in Figure 15 and run generation with temperature 0.3 and no top-k sampling.
 You are responsible for rewriting prompts for a computer program that generates images.
@@
 Figure 15: System prompt for Llava-v1.6-34b when generating iteratively more difficult prompts.
-MT2IE Adaptive Prompting Prompt Generation (Section section 6)
-For adaptive prompt generation in Section section 6, we use the system prompts shown in Figures figure 16, figure 17, figure 18 and run generation with temperature 0.3 and no top-k sampling.
+MT2IE Adaptive Prompting Prompt Generation (Section 6)
+For adaptive prompt generation in Section 6, we use the system prompts shown in Figures 16, 17, 18 and run generation with temperature 0.3 and no top-k sampling.
 Score $[0.0,0.2]$:
@@
 MT2IE: 12 Prompt Iterations (48 Prompts) $0.7025$
-MT2IEgenerates a configurable number of prompts: in the Iterative Difficulty setting discussed in Section section 5 pronpts are progressively more difficulty in each iteration, whereas in the Adaptive Prompting setting discussed in Section section 6 each iteration’s prompt is rewritten by the MLLM based on the previous prompt’s alignment score.
-Rankings of 8 T2I models for prompt iterations={2, 5, 10, 12} for Iterative Difficulty are shown in Figure figure 19 and Adaptive Prompting in Figure figure 20.
-Additionally we calculate Kendall’s $\tau$ rank correlation coefficient for all T2I model rankings, using the ranking obtained from GenAIBench as the ground truth ranking, shown in Table table 5.
+MT2IEgenerates a configurable number of prompts: in the Iterative Difficulty setting discussed in Section 5 pronpts are progressively more difficulty in each iteration, whereas in the Adaptive Prompting setting discussed in Section 6 each iteration’s prompt is rewritten by the MLLM based on the previous prompt’s alignment score.
+Rankings of 8 T2I models for prompt iterations={2, 5, 10, 12} for Iterative Difficulty are shown in Figure 19 and Adaptive Prompting in Figure 20.
+Additionally we calculate Kendall’s $\tau$ rank correlation coefficient for all T2I model rankings, using the ranking obtained from GenAIBench as the ground truth ranking, shown in Table 5.
 We found that using 5 prompt iterations, i.e.,  20 total prompts, with MT2IEyields strongly-correlated model scores with the GenAIBench benchmark which contains 1600 prompts.
@@
 Despite this, Llava-Gen TIFA achieves higher correlations with human judgment than TIFAv1.0, supporting our claims that MLLM-generated T2I benchmarks are more efficient.
-The distribution of question counts per prompt are shown in Figure figure 21.
+The distribution of question counts per prompt are shown in Figure 21.
 Similarity of question elements
@@
 Across all prompts, TIFAv1.0 has an average of 4.68 elements per prompt and Llava-Gen TIFA has 4.13 elements per prompt, indicating an extremely similar number of elements identified for each prompt.
-For prompts with mismatched elements, the average number of mismatched elements is only 1.13, and we show counts of prompts with mismatched elements in Figure figure 22.
+For prompts with mismatched elements, the average number of mismatched elements is only 1.13, and we show counts of prompts with mismatched elements in Figure 22.
 The average BLEU (Papineni et al., 2002) score between TIFAv1.0 and Llava-Gen TIFA questions about the same elements is 0.65, which indicates a high level of similarity between questions covering the same concepts.